### v3

# Material-informed Gaussian Splatting for 3D World Reconstruction in a Digital Twin

Abstract
3D reconstruction for Digital Twins often relies on LiDAR-based methods, which provide accurate geometry but lack the semantics and textures naturally captured by cameras. Traditional LiDAR-camera fusion approaches require complex calibration and still struggle with certain materials like glass, which are visible in images but poorly represented in point clouds. We propose a camera-only pipeline that reconstructs scenes using 3D Gaussian Splatting from multi-view images, extracts semantic material masks via vision models, converts Gaussian representations to mesh surfaces with projected material labels, and assigns physics-based material properties for accurate sensor simulation in modern graphics engines and simulators. This approach combines photorealistic reconstruction with physics-based material assignment, providing sensor simulation fidelity comparable to LiDAR-camera fusion while eliminating hardware complexity and calibration requirements. We validate our camera-only method using an internal dataset from an instrumented test vehicle, leveraging LiDAR as ground truth for reflectivity validation alongside image similarity metrics.
Figure 1: Overview of our camera-only reconstruction pipeline. From RGB images, we: (A) extract semantic material masks, (B) reconstruct the 3D scene, (C) project material labels onto mesh surfaces, (D) assign physics-based materials, and (E) validate through sensor simulation.
I Introduction
Digital Twins advance sensor technologies by enabling safe validation of high-risk scenarios and precise environmental control prior to real-world deployment. Modern Advanced Driver Assistance Systems (ADAS) and Artificial Intelligence (AI) systems rely on sophisticated algorithms and complex multimodal sensor setups, often including LiDAR sensors and cameras. Achieving accurate translation of real-world scenarios to simulation—where virtual sensors exhibit identical behavior to their physical counterparts—requires 3D world reconstruction that captures spatial and geometric features to ensure realistic simulation and synthetic sensor data consistent with reality.
Previous LiDAR-based approaches Zaalberg2024 ; Manivasagam achieve accurate geometry but lack camera-captured texture information. Moreover, LiDAR struggles with transparent and reflective materials like glass and metal, readily captured in camera images. While incorporating material properties into LiDAR-based reconstruction has been shown to improve sensor simulation accuracy muckenhuber2020automotive , such approaches still require LiDAR hardware and do not address texture acquisition. Combining LiDAR with cameras to capture both geometry and appearance requires complex sensor synchronization and calibration procedures, yet still produces limited texture quality under sparse-view conditions, as traditional reconstruction methods struggle to generate photorealistic results.
Recent advancements in novel view synthesis have introduced 3D Gaussian Splatting kerbl20233dgaussiansplattingrealtime , an efficient technique providing fast, differentiable rendering and high-quality reconstructions from sparse input views. Recent works like MILo guédon2025milomeshintheloopgaussiansplatting extract geometric surfaces from Gaussian Splatting using only multi-view images, enabling hybrid representations combining photorealistic visualization with physics-based simulation in graphics engines. These camera-only methods provide an alternative to LiDAR hardware, avoiding intricate sensor synchronization and demanding calibration procedures while delivering superior texture quality.
Building on these observations, we propose a camera-only pipeline that combines photorealistic Gaussian Splatting with physics-based material assignment (Fig. 1). This approach enables accurate sensor simulation by integrating semantic material information into visually realistic reconstructions.
Our main contributions are:
  • (1) An automated method for projecting 2D semantic material masks onto 3D mesh surfaces through Gaussian Splatting-based reconstruction, enabling accurate physics-based LiDAR reflectivity simulation.
  • (2) A modular, camera-only pipeline integrating Gaussian Splatting for photorealistic geometry reconstruction with automated mesh-based material assignment.
  • (3) Comprehensive evaluation demonstrating sensor simulation accuracy comparable to LiDAR-camera fusion, validated through LiDAR reflectivity and rendering quality analysis.
While our pipeline requires only camera input, we validate it using LiDAR ground truth from an instrumented test vehicle.
II Related Work
II-A Novel View Synthesis
Traditional 3D reconstruction methods use explicit representations. LiDAR-based methods huang2023neural excel at geometric accuracy but struggle with fine visual details. Camera-based alternatives Schonberger2016 offer richer visual information but produce noisier geometry and face challenges with transparent and reflective surfaces remondino2023critical . Neural Radiance Fields (NeRF) mildenhall2020nerfrepresentingscenesneural enabled photorealistic novel view synthesis from camera-only input but suffers from high computational overhead and lacks discrete surfaces needed for graphics engines.
3D Gaussian Splatting kerbl20233dgaussiansplattingrealtime addressed these limitations through an explicit point-based representation enabling real-time rendering with photorealistic quality. Subsequent refinements have improved geometric accuracy, transparency handling, and shading Huang_2024 ; zhang2024quadraticgaussiansplattingefficient ; wu20253dgut . However, these methods remain primarily oriented toward small, object-centric captures and struggle with large-scale outdoor scenes.
II-B Large-Scale Scene Reconstruction
To scale Gaussian Splatting to large-scale outdoor environments, recent methods have adopted scene partitioning strategies tancik2022blocknerfscalablelargescene , dividing scenes into spatially consistent blocks for parallel reconstruction liu2025citygaussianv2efficientgeometricallyaccurate ; h3dgs . Many employ Level-of-Detail (LoD) techniques, allocating higher resolution to nearby regions to improve rendering speed. Alternative strategies include temporal segmentation cui2024streetsurfgsscalableurbanstreet and distance-based decomposition shi2024dhgsdecoupledhybridgaussian . However, these methods often require manual parameter tuning and lack robust surface representations needed for simulation engines.
II-C Surface Extraction from Gaussian Primitives
While Gaussian Splatting enables high-quality rendering, simulation engines require explicit mesh representations. Methods such as guedon2023sugar use Poisson reconstruction, while Signed Distance Field (SDF)-based approaches chen2025neusgneuralimplicitsurface require dense multi-view data. Recent work guédon2025milomeshintheloopgaussiansplatting integrates mesh extraction directly into training via Delaunay triangulation, enabling joint optimization of appearance and geometry. However, these methods are primarily designed for controlled, object-centric scenarios and lack robustness for large-scale outdoor environments.
II-D Material Extraction
Beyond extracting accurate geometry, material properties are essential for realistic simulation. Image-based methods classify materials through semantic segmentation cai2022rgb . Multi-modal approaches Liang_2022_CVPR and large-scale datasets upchurch2022densematerialsegmentationdataset have improved accuracy. While LiDAR-based methods muckenhuber2020automotive ; viswanath2024reflectivityneedadvancinglidar measure physical properties directly, camera-only approaches capture materials that LiDAR struggles with, such as transparent and reflective surfaces. However, transferring 2D material segmentation to 3D reconstructed surfaces remains challenging for camera-only Gaussian Splatting pipelines, where material labels must be accurately projected from multi-view observations onto extracted mesh geometry. Our work addresses this gap by introducing an automated projection method that enables physics-based sensor simulation.
Figure 2: Shape-aware material refinement. From left to right: input RGB image, RMSNet predictions, FastSAM object boundaries, and final result after majority voting. The combination produces consistent material labels with sharp edges. Example from our internal dataset.
II-E Physics-Based Materials for Rendering
Once material labels are assigned to 3D surfaces, they must be represented in a format compatible with physics-based rendering engines. Modern graphics engines simulate surface interactions through the Bidirectional Reflectance Distribution Function (BRDF) murat2018survey . Recent work integrates material properties directly into 3D Gaussians ye2025largematerialgaussianmodel ; yao2025reflectivegaussiansplatting focusing on rendering, while mesh-compatible approaches Xiong_2025_CVPR enable seamless integration with standard simulation engines. However, bridging the gap between semantic material labels from 2D segmentation and physically accurate material parameters for physics-based sensor simulation in Digital Twin environments remains an open challenge.
III Method
Our pipeline reconstructs large-scale outdoor scenes from multi-view RGB images and assigns physics-based materials for accurate sensor simulation. The approach is decoupled, separating geometric reconstruction from material assignment, and employs a hybrid representation, combining photorealistic Gaussian Splatting with explicit mesh geometry. As illustrated in Fig. 1, the pipeline consists of five stages: monocular material extraction (Sec. III-A), Gaussian Splatting reconstruction (Sec. III-B), per-pixel material projection from 2D to 3D (Sec. III-C), physics-based material assignment (Sec. III-D), and simulation validation (Sec. III-E).
III-A Monocular Material Extraction
We extract per-pixel material labels from RGB images using a two-stage approach: texture-based material segmentation followed by shape-aware refinement. For the first stage, we adopt RMSNet cai2022rgb due to its strong performance on autonomous driving datasets and its ability to operate on RGB images alone, without requiring LiDAR or infrared sensors. It distinguishes classes such as asphalt, concrete, glass, and metal.
Material regions often exhibit fragmented distributions lacking prominent shape cues. We employ FastSAM zhao2023fastsegment to identify instances, producing pixel-wise boundaries that we post-process to remove overlaps before overlaying with RMSNet predictions.
For each region, we apply majority voting: counting material class labels within its boundary and assigning the most frequent label to all enclosed pixels. This strategy ensures consistent material assignment while preserving sharp boundaries, as illustrated in Fig. 2. The result is a set of n shape-aware material masks, where each pixel inherits its class according to the FastSAM-derived segmentation.
III-B Large-Scale Gaussian Splatting
Our 3D reconstruction combines photorealistic rendering with explicit geometry through a hybrid approach. We first apply COLMAP Schonberger2016 to generate a sparse point cloud from multi-view RGB images, providing GPS-derived camera poses, pre-calibrated intrinsics, and extrinsics as priors. We mask the moving ego-vehicle in all input images to avoid artifacts.
From the COLMAP output, we train two complementary models. For photorealistic visualization, we use H3DGS h3dgs , selected for its superior visual quality on autonomous driving datasets over alternatives like CityGaussianV2 liu2025citygaussianv2efficientgeometricallyaccurate . H3DGS constructs a global scaffold and subdivides the scene into chunks. However, H3DGS’s hierarchical LoD representation uses a proprietary .hier format, and its anti-aliasing mechanism yu2023mip is not supported by standard 3DGS rendering tools that expect the conventional .ply format. To maximize compatibility with existing 3DGS viewers and downstream processing tools, we therefore disable these features and perform a custom merging of the chunks.
For mesh geometry and semantic labeling, we employ MiLO guédon2025milomeshintheloopgaussiansplatting , which provides geometrically accurate mesh extraction required for physics-based simulation. Trained on the same COLMAP sparse point cloud, MiLO produces both a 3D Gaussian model and an explicit mesh. The MiLO Gaussian model initializes our semantic Gaussian Splatting for material projection (Sec. III-C), while the mesh provides the geometric basis for physics-based simulation (Sec. III-E). Both H3DGS and MiLO are trained with monocular depth supervision depth_anything_v2 following their respective default configurations, and both reconstructions share the same coordinate system, ensuring alignment between appearance and geometry.
Figure 3: Qualitative comparison of novel view rendering. From left to right: ground truth RGB, our adapted H3DGS model, 3DGS, CityGaussianV2, and 3DGUT. Our model achieves competitive visual quality comparable to state-of-the-art baselines.
III-C Per-Pixel Material Projection
This module projects the 2D material masks from Sec. III-A onto the MiLO Gaussian representation, then transfers the labels to the mesh. Unlike previous pipelines relying on camera-to-LiDAR transformations, our Gaussian model is natively aligned in world coordinates. We employ differentiable rasterization to render Gaussians to 2D, maintaining correspondence between 3D Gaussian positions and 2D pixel coordinates. When multiple Gaussians project to the same pixel, depth sorting during alpha blending assigns the label of the closest Gaussian.
Inconsistencies across views, such as static objects receiving different labels, are resolved using SegAnyGS cen2025segment3dgaussians , which enforces consistent assignment of our material masks to the Gaussians. We apply majority voting across overlaid masks to ensure label consistency. Each material-labeled Gaussian is assigned to its nearest mesh triangle using K-Nearest Neighbors (KNN), and the triangle inherits the Gaussian’s label, producing a per-triangle material mesh as the final output.
III-D Physics-Based Material Assignment
We assign physics-based material properties to the mesh using the Principled BSDF shader following the Disney BSDF standard burley2012physically , ensuring compatibility with graphics engines such as Blender, Unreal Engine, and simulation platforms like Simcenter Prescan. The Principled BSDF is parameterized by physically meaningful properties including base color, specular, metallic factor, roughness, opacity, and surface normals.
From the 20 material classes detected by RMSNet (Sec. III-A), we selected a subset commonly encountered in urban driving environments: glass, brick and ceramics, concrete, asphalt, vegetation (e.g., grass and leaves), metals, plastics, gravel, tree trunks, and rubber. For each category, PBR textures were sourced from the Matsynth dataset Vecchio_2024 , which provides over 4,000 physically-based materials across 14 categories, and matched to Prescan’s laboratory-tested material database to ensure validated reflectivity properties for LiDAR simulation. Within each category, textures were balanced across key parameters (base color, roughness, metallic factor, clear coat). For each material class from Sec. III-A, we apply the selected PBR textures to the corresponding mesh triangles, producing a fully textured mesh directly usable by standard graphics engines and physics-based sensor simulation frameworks.
III-E Simulation Validation
To validate our material-informed reconstruction pipeline, we integrate the reconstructed scenes into Simcenter Prescan to recreate realistic traffic scenarios with physics-based sensor simulation. The material-labeled mesh from Sec. III-D is imported into Prescan, where laboratory-measured PBR material properties are assigned to each triangle based on its material label. This enables accurate interaction with physics-based sensors, particularly LiDAR, where surface reflectivity depends on material composition.
We replicate the original sensor configuration from our instrumented test vehicle, including LiDAR specifications and mounting positions. The ego-vehicle trajectory is extracted and imported via the Simcenter Prescan MATLAB API, ensuring the virtual sensor follows the exact path and poses as during real-world data capture. The simulation generates synthetic LiDAR point clouds with material-specific reflectivity values for each point.
We validate our camera-only reconstruction by comparing these simulated LiDAR responses against real-world LiDAR measurements from the instrumented test vehicle. This demonstrates that our pipeline produces 3D models suitable for physics-based sensor simulation, providing a practical alternative to traditional LiDAR-based reconstruction workflows while achieving comparable sensor simulation accuracy.
TABLE I: Image similarity metrics (PSNR $\uparrow$, SSIM $\uparrow$, LPIPS $\downarrow$) evaluated on novel views for all baselines.
Adapted H3DGS (Ours) 3DGS CityGaussianV2 3DGUT
[…]
Scene 1 16.88 0.299 0.451 19.13 0.476 0.438 20.36 0.486 0.480 19.97 0.465 0.517
Scene 2 17.86 0.358 0.432 19.81 0.525 0.474 20.79 0.526 0.522 21.16 0.505 0.512
Scene 3 18.35 0.375 0.437 19.48 0.515 0.492 20.42 0.524 0.540 21.10 0.502 0.557
Scene 4 19.03 0.504 0.439 20.88 0.659 0.450 21.15 0.676 0.483 21.85 0.666 0.526
Scene 5 19.21 0.503 0.427 21.04 0.650 0.446 21.63 0.690 0.499 22.04 0.662 0.520
Average 18.27 0.41 0.44 20.07 0.57 0.46 20.87 0.58 0.50 21.22 0.56 0.53
  • Best performance highlighted in red, second-best in yellow.
IV Experiments
We evaluate our camera-only reconstruction pipeline using our internal dataset. To assess whether camera-only reconstruction achieves sensor simulation accuracy comparable to LiDAR-based methods, we establish a LiDAR-camera fusion baseline that combines LiDAR-based mesh reconstruction (NKSR huang2023neural ) with camera-derived material labels. This baseline provides geometric accuracy through direct depth measurements while incorporating semantic material information from camera images.
We conduct two complementary evaluations with different baselines: (1) visual quality, where we compare our Gaussian Splatting reconstruction against state-of-the-art camera-only methods (3DGS kerbl20233dgaussiansplattingrealtime , 3DGUT wu20253dgut , CityGaussianV2 liu2025citygaussianv2efficientgeometricallyaccurate ) using image similarity metrics (PSNR, SSIM, LPIPS), and (2) sensor simulation accuracy, where we compare our camera-only approach against the LiDAR-camera fusion baseline by evaluating synthetic LiDAR reflectivity against real-world ground truth.
IV-A Dataset
We collected data using an instrumented test vehicle equipped with an Ouster OS1-128 LiDAR ousterOS1Specs (128 channels, 20 Hz, 100 Hz IMU), six surround-view cameras with Sony IMX490 sensors (2880×1860 resolution, one rear with fisheye lens for 360° coverage), and a Septentrio AsteRx GNSS/RTK system (20 Hz). The LiDAR provides hardware triggers to synchronize all cameras at 20 Hz. Five static road scenes were captured near our company facilities in Leuven, Belgium, each approximately eight seconds at 20-30 km/h. Data collection on company-controlled roads avoided GDPR compliance issues and ensured controlled validation conditions.
We focus on static scene reconstruction as this aligns with Digital Twin simulation workflows. Modern simulation platforms (e.g., Simcenter Prescan, CARLA) provide realistic dynamic actor models with controllable behaviors, making the challenge the creation of photorealistic static environments (roads, buildings, vegetation, infrastructure). Our method automates this process, providing realistic backgrounds for simulation scenarios. Dynamic actors in our recordings are masked out during reconstruction. While our detection pipeline enables trajectory replication for re-insertion as simulation assets, this falls outside the scope of this paper.
Scenes primarily consist of company parking lots and nearby access roads, capturing typical automotive testing environments with diverse material challenges including asphalt, concrete, glass facades, metallic surfaces, extensive vegetation (trees, hedges, grass), road markings, and signage. Recordings under varying weather conditions (clear to partly cloudy) capture natural lighting variations. We reserve 50 novel viewpoints per scene for evaluation.
IV-B Baseline
Visual quality: We compare our Gaussian Splatting reconstruction against three state-of-the-art camera-only methods: 3DGS kerbl20233dgaussiansplattingrealtime , 3DGUT wu20253dgut , and CityGaussianV2 liu2025citygaussianv2efficientgeometricallyaccurate . All methods are trained on the same five scenes and evaluated on 50 uniformly sampled novel viewpoints per scene using PSNR, SSIM, and LPIPS.
Sensor simulation accuracy: We compare our camera-only reconstruction against a LiDAR-camera fusion baseline by simulating both environments in Simcenter Prescan with identical PBR material properties and ego-vehicle trajectories. The LiDAR-based baseline uses NKSR huang2023neural for mesh reconstruction from LiDAR point clouds, combined with the same camera-derived material labels as our camera-only approach. This controlled setup minimizes external variables, allowing direct comparison of geometric reconstruction quality and its impact on reflectivity accuracy.
Figure 4: From left to right: real-world LiDAR ground truth from our instrumented vehicle, LiDAR-camera fusion baseline simulated in Prescan, and our camera-only reconstruction simulated in Prescan.
IV-C Implementation Details
Scene reconstruction: We train all Gaussian Splatting methods using their default hyperparameters as specified in their respective repositories. For both H3DGS and MiLO, we enable monocular depth supervision using Depth Anything V2 depth_anything_v2 . Material assignment follows the pipeline described in Sec. III-AIII-D, projecting semantic labels onto the reconstructed geometry and mapping them to PBR material properties.
LiDAR simulation: We configure the Simcenter Prescan LiDAR sensor according to the specifications of the Ouster OS1-128 ousterOS1Specs used in our dataset. The physics-based sensor returns both Cartesian coordinates of detected points and a power value representing reflected signal strength. We normalize this power output for range and incidence angle to obtain accurate surface reflectivity values, following established calibration procedures hermidas_phillips_physlidarpowercal .
Hardware: All experiments were conducted on a Dell Precision 7680 equipped with 64GB RAM and an NVIDIA RTX A4000 ADA Laptop GPU. Sensor validation was performed using Simcenter Prescan 2411, with scenario control via the Prescan MATLAB API (R2023b).
IV-D Metrics
We employ complementary metrics for visual quality and sensor simulation accuracy:
Visual quality: PSNR gonzalez2008digital (pixel-level fidelity), SSIM 1284395 (structural similarity), and LPIPS zhang2018unreasonableeffectivenessdeepfeatures (perceptual quality). Higher is better for PSNR and SSIM; lower is better for LPIPS.
LiDAR simulation accuracy: Mean Absolute Error (MAE) and median error quantify reflectivity prediction error between synthetic and real-world measurements, with median providing robustness against outliers.
IV-E Visual Quality Evaluation
We evaluate our adapted H3DGS model against three state-of-the-art Gaussian Splatting methods by comparing novel view renderings across all five scenes.
Qualitative analysis. Figure 3 demonstrates that our adapted model achieves competitive visual quality with minimal deviation from ground truth. The model effectively fills masked ego-vehicle regions and produces cleaner reconstructions by filtering sensor noise in the ground truth. However, limitations emerge when reconstructing highly reflective and transparent surfaces, occasionally resulting in misalignments with ground truth. Rendering quality is highest for front-facing cameras, while side cameras exhibit reduced quality due to motion blur from the vehicle’s movement. Scenes with dense object distributions and occlusions also reduce reconstruction quality.
Quantitative analysis. Table I presents PSNR, SSIM, and LPIPS metrics across all methods and scenes. Our adapted H3DGS model is used solely for photorealistic visualization, while MiLO provides geometry for simulation (Sec. III-B). The model achieves lower PSNR (18.27 vs. 20.07–21.22) and SSIM (0.41 vs. 0.56–0.58) compared to baselines, but the best LPIPS score (0.44 vs. 0.46–0.53), outperforming all methods on four out of five scenes. Since LPIPS correlates strongly with human perceptual similarity zhang2018unreasonableeffectivenessdeepfeatures , this suggests our reconstruction maintains perceptual quality despite lower pixel-level accuracy.
The degradation stems from disabling H3DGS’s hierarchical optimization and anti-aliasing mechanism yu2023mip . As noted by the H3DGS authors, hierarchies operating at different scales require correct anti-aliasing h3dgs . Anti-aliasing alone improves 3DGS quality by +4.56 dB PSNR and +0.070 SSIM yu2023mip ; our observed degradation (-1.80 dB PSNR, -0.15 SSIM) is consistent with these modifications. The superior LPIPS score confirms these changes preserve perceptual quality while ensuring compatibility with standard rendering tools.
IV-F LiDAR Simulation Accuracy
We evaluate our camera-only Gaussian Splatting reconstruction against the LiDAR-camera fusion baseline by measuring reflectivity error relative to real-world ground truth.
TABLE II: LiDAR reflectivity prediction error (lower is better). Reflectivity values normalized to 0–255 range.
Camera-only (Ours) LiDAR-based
$downarrow MAE\downarrow Median\downarrow$
Scene 1 10.91 7.16 11.73 7.84
Scene 2 9.37 5.54 10.17 6.70
Scene 3 11.03 6.91 10.47 6.87
Scene 4 9.32 5.77 9.41 6.02
Scene 5 8.98 6.71 8.76 6.36
Average 10.05 6.48 10.14 6.78
  • MAE: Mean Absolute Error. Our camera-only method achieves comparable accuracy to the LiDAR-based baseline.
Qualitative analysis. Figure 4 presents a visual comparison of LiDAR reflectivity predictions between our camera-only method, the LiDAR-camera fusion baseline, and real-world ground truth. Both methods consistently overestimate vegetation reflectivity due to: (1) temporal inconsistency from wind-induced vegetation movement violating static scene assumptions, (2) high intra-class material variance (leaves, bark, grass share one label but have different reflectivities), (3) idealized PBR materials not capturing weathered outdoor conditions, and (4) geometric complexity affecting surface normal estimation. These limitations impact both reconstruction approaches equally, indicating bottlenecks in the static scene assumption and material parameterization rather than reconstruction modality. Despite higher geometric noise in the camera-only method, both approaches produce comparable reflectivity patterns, validating our pipeline for sensor simulation.
Figure 5: Distribution of absolute reflectivity errors for camera-only (ours) and LiDAR-based reconstruction methods. Our camera-only approach achieves comparable median error with slightly higher variability.
Quantitative analysis. Table II and Figure 5 present reflectivity error metrics across all five scenes. Our camera-only method achieves comparable reflectivity accuracy to the LiDAR-camera fusion baseline, with mean absolute error (MAE: 10.05 vs. 10.14) and median error (6.48 vs. 6.78) showing minimal differences between the two approaches. This demonstrates that camera-only reconstruction can achieve sensor simulation accuracy comparable to LiDAR-based methods without requiring LiDAR hardware or complex sensor synchronization. The higher within-scene variability in camera-only results (visible in Figure 5) reflects the increased geometric noise inherent to MiLO’s camera-based mesh extraction, which lacks the direct depth measurements provided by LiDAR. However, the comparable MAE and median errors demonstrate that this geometric uncertainty does not significantly degrade reflectivity simulation accuracy.
IV-G Ablation Study
We evaluate the impact of incorporating segmentation masks to improve material classification accuracy. As described in Sec. III-A, we augment RMSNet’s material predictions with shape cues from segmentation models.
Experimental setup. We quantitatively assess segmentation quality using the MCubes multi-modal dataset Liang_2022_CVPR , which contains 500 material-annotated images. We compare three configurations: (1) RMSNet alone, (2) RMSNet + FastSAM, and (3) RMSNet + SAM2 ravi2024sam2segmentimages .
TABLE III: Material segmentation performance on MCubes dataset Liang_2022_CVPR .
Method Accuracy mIoU
RMSNet 0.63 0.28
RMSNet + FastSAM (Ours) 0.65 0.29
RMSNet + SAM2 0.67 0.31
Results. Table III shows that adding segmentation masks improves both accuracy (from 0.63 to 0.65–0.67) and mIoU (from 0.28 to 0.29–0.31) over RMSNet alone. While SAM2 achieves the highest scores (0.67 accuracy, 0.31 mIoU), we select FastSAM (0.65 accuracy, 0.29 mIoU) due to its significantly lower computational cost and faster inference speed—critical for processing large-scale multi-view datasets. Although SAM2 offers superior segmentation quality, the 2% accuracy improvement does not justify the increased computational overhead for our application.
Analysis. The 3.2% accuracy improvement (0.63 to 0.65) demonstrates that shape cues effectively enhance material classification consistency. However, segmentation masks can occasionally reduce per-pixel accuracy when incorrect predictions propagate across larger regions, potentially overwriting fine details such as road markings. The 2% accuracy gap between FastSAM and SAM2 (0.65 vs. 0.67) represents an acceptable tradeoff for the practical benefits of automatic mask extraction in large-scale scenarios.
V Conclusion
We present a camera-only pipeline that combines photorealistic Gaussian Splatting with physics-based material assignment for Digital Twin reconstruction. Our approach leverages Gaussian Splatting to bridge 2D camera observations to comprehensive 3D scene representations, including photorealistic visuals, semantic information, mesh geometry, and material properties for sensor simulation. Through comprehensive evaluation on real-world urban driving scenes using LiDAR ground truth for validation, we demonstrate that our material projection approach achieves reflectivity prediction accuracy comparable to LiDAR-based methods while maintaining photorealistic rendering quality, offering a practical alternative for ADAS development when LiDAR is unavailable or impractical.
Future work. Future directions include improving geometric reconstruction quality from camera inputs, extending to dynamic scenes, and exploring learned material representations to reduce dependency on predefined PBR databases. Specific improvements for vegetation modeling could include: (1) environment-specific material calibration capturing weathering and seasonal effects, (2) hierarchical material classification distinguishing foliage types, bark, and ground cover, and (3) refined geometric alignment between multi-modal reconstructions to improve surface normal accuracy for reflectivity calculations. We have developed an internal tool to convert our data to the nuScenes format, enabling future benchmarking against public datasets to demonstrate broader generalization across diverse urban environments.